# FC-KAN: Function Combinations in Kolmogorov-Arnold Networks

### v3

@@ 2.1 KART and KAN
In 1957, Kolmogorov provided a proof to Hilbert’s 13th problem by showing that any multivariate continuous function can be expressed as a combination of single-variable functions and additions, a concept known as KART [4, 16]. This theorem has been utilized in many studies to develop neural networks [17, 18, 19, 20, 21, 22]. However, there is an ongoing debate about the applicability of KART in neural network design. Girosi and Poggio [23] argued that KART’s relevance to neural networks is questionable because the inner function $\phi_{q,p}$ in Equation 1 may be highly non-smooth [24], which could hinder $f$ from being smooth—a key attribute for generalization and noise resistance in neural networks. Conversely, Kůrková [25] contended that KART is applicable to neural networks, showing that linear combinations of affine functions can effectively approximate all single-variable functions using certain sigmoidal functions.
@@ 3.1 Kolmogorov-Arnold Representation Theorem
A KAN is based on KART, which states that any continuous multivariate function $f$ defined on a bounded domain can be represented as a finite combination of continuous single-variable functions and their additions [60, 61]. For a set of variables $\mathbf{x}={x_{1},x_{2},\ldots,x_{n}}$, where $n$ is the number of variables, the multivariate continuous function $f(\mathbf{x})$ is expressed as:
@@
which has two types of summations: the outer sum and the inner sum. The outer sum, $\sum_{q=1}^{2n+1}$, aggregates $12n+$ 1 terms of $\Phi_{q}$ ($\mathbb{R}\to\mathbb{R}$). The inner sum, on the other hand, aggregates $n$ terms for each $q$, where each term $\phi_{q,p}$ ($\phi_{q,p}\colon[0,1]\to\mathbb{R}$) denotes a continuous function of a single variable $x_{p}$.
@@ 3.2 Design of KANs
Remind an MLP that consists of affine transformations and non-linear functions. Starting with an input $\mathbf{x}$, the network processes it through a series of weight matrices across layers (from layer 00 to layer $L-1$) and applies the non-linear activation function $\sigma$ to produce the final output.
@@
Inspired by KART, Liu et al. [1] developed KANs but recommended extending the approach to incorporate greater network widths and depths. To address this, appropriate functions $\Phi_{q}$ and $\phi_{q,p}$ in Equation 1 need to be identified. A typical KAN network with $L$ layers processes the input $\mathbf{x}$ to produce the output as follows:
@@ 3.3 Implementation of the Current KANs
where $b(x)$ equals to $silu(x)$ as in Equation 7 and $spline(x)$ is expressed as a linear combination of B-splines $B_{i}$s and control points (coefficients) $c_{i}$s as in Equation 8. Each activation function is activated with $w_{s}=1$ and $spline(x)\approx 0$, while $w_{b}$ is initialized by using Xavier initialization.
Figure 1 shows the architecture of KAN(2,3,1), which includes 2 input nodes, 3 hidden nodes, and 1 output node. The output of each node is derived from the sum of individual functions $\phi$, called "edges." The diagram also explains the computation of the inner function $\phi$ using control points (coefficients) and B-splines. The number of B-splines is determined by adding the grid size $G$ and the spline order $k$, resulting in $G+k=3+3=6$, corresponding to the range of $i$ from 0 to 5.
@@
[…]
$=\sum_{i=1}^{N}w_{i}\left(1-\left(\tanh\left(\frac{||x-c_{i}||}{h}\right)\right)^{2}\right)$
@@
For the simulation of function plots, we present them in Figure 2, where they are fitted to pass through four selected points. These plots provide a visual representation of the different function types and how they interact with the data points, helping to illustrate their shapes, which are not intended for pre-validating model performance trained on these functions. The functions analyzed include B-spline (3rd-degree), DoG, GRBF, RSWAF, and BSRBF (a combination of B-spline and GRBF). Among these, only the B-spline perfectly passes through all points, while DoG, BSRBF, and GRBF intersect with one point each, and RSWAF passes through none. It is important to note that a function passing through all specified points does not guarantee strong model performance in neural networks, as issues such as overfitting, poor generalization, and sensitivity to outliers may arise.
@@ 3.4 FC-KAN
Figure 4: Various data combinations are performed using element-wise operations (additions $+$ and multiplications $\odot$) over two given outputs. The outputs always have the same data dimensions as the inputs.
@@
After that, we combine function outputs to obtain the element-wise output, the concatenated output, and the linearized output, as shown in Figure 3. The element-wise output is formed by using element-wise operations, including the sum output (Equation 17a), the element-wise product output (Equation 17b), the sum and element-wise product output (Equation 17c), the quadratic and cubic function outputs (Equation 17d and Equation 17e). The concatenated output is formed by simply concatenating the function outputs together as in Equation 17f. The linearized output is formed by passing the concatenated output through a linear transformation (with matrix multiplication by $W$ and addition of bias $b$) as in Equation 17g. The element-wise output and the linearized output will maintain the same data dimension as each element $\mathbf{o_{i}}$ in $O$, except in the concatenated output, where the data size will be $\mathbf{o_{i}}$ multiplied by the number of function outputs being combined.
@@
Figure 4 presents data combinations over two given function outputs $\mathbf{o_{1}}$ and $\mathbf{o_{2}}$ using element-wise operations. The results are 5 combined outputs: sum, sum + prod, prod, quad, and cubic. Output combinations can utilize higher-degree functions, but these may significantly increase computational complexity, especially in matrix multiplication. Additionally, using more functions results in a larger number of outputs, which can further complicate data combination calculations. To manage this complexity, we prefer to restrict output combinations to quadratic functions involving up to two outputs. For instance, to combine DoG and B-splines at the output, we can use the following quadratic function formula:
@@ 4.1 Training Configuration
As shown in Table 1, all models contain a network structure of (784, 64, 10), comprising 784 input neurons, 64 hidden neurons, and 10 output neurons corresponding to the 10 output classes (0-9). Due to the function combinations, FC-KAN has the highest number of parameters, while the MLP has the fewest because it only contains linear transformations over data between layers. The models were trained with 25 epochs on MNIST and 35 epochs on Fashion-MNIST. For KAN models, we use grid_size=5, spline_order=3, and num_grids=8. Other hyperparameters are the same in all models, including batch_size=64, learning_rate=1e-3, weight_decay=1e-4, gamma=0.8, optimize=AdamW, and loss=CrossEntropy.
@@ 4.2 Model Performance
Figure 5 shows the training losses, represented on a logarithmic scale, for MLP and KAN models on the MNIST and Fashion-MNIST datasets. The loss performance of each model was evaluated based on an independent training run. FC-KAN consistently achieves the lowest training losses across both datasets, followed by BSRBF-KAN due to its fast convergence feature. In contrast, FasterKAN records the highest training loss on MNIST, while MLP performs similarly on Fashion-MNIST.
In general, FC-KAN models outperformed others on MNIST and Fashion-MNIST but require more training time due to the quadratic function representation for output combination, as shown in Section 4.2. This trade-off between training time and model performance is considered reasonable. The best-performing models are FC-KAN (BS+BASE) and FC-KAN (DoG+BS), which achieved validation accuracies of 97.93% on MNIST and 89.99% on Fashion-MNIST, respectively. When calculating the metric values for both datasets, FC-KAN (DoG+BS) slightly surpassed FC-KAN (BS+BASE) and outperformed the other models. However, FC-KAN (BS+BASE) models take between 9.25% and 11.55% less training time.
@@ 4.3 Comparison of Combination Methods
From the results in Section 4.3, the quadratic function best represents the combination output and outperforms other combinations, although its models require more training time. It is clear that the outputs combined by element-wise operations consistently outperform other methods, demonstrating superior accuracy. This indicates that element-wise combinations are more effective in capturing and integrating relevant features from the data, leading to better performance. Meanwhile, the output combination by concatenation shows the worst results. The linear transformation applied to the concatenated output outperforms concatenation alone, but still achieves only average performance. Similarly, the maximum, minimum, and average outputs do not deliver superior results.
@@ 4.4 Misclassification Analysis
Figure 6: Heatmaps of the misclassified images in the validation set by models over MNIST and Fashion-MNIST.
@@ 4.5 Model Performance with Limited Data
Figure 7: The validation accuracy values of the models across various data subsets.
@@
Figure 7 illustrates the performance of the FC-KAN models across different training subsets and function combinations. In the MNIST dataset, the RBF+BASE combination demonstrates superior performance with 1%, 5%, and 10% of the data; however, the highest performance on the full training data is recorded for the BS+BASE combination. This observation underscores the challenges of predicting the optimal function combination when working with limited training data. In the Fashion-MNIST dataset, DoG+BS achieves the best performance with 1% of the data, while DoG+RBF excels with both 5% and 10%. These results suggest that DoG+BS and DoG+RBF may perform well with the full training dataset. Indeed, DoG+BS consistently outperforms other combinations in full training, with DoG+RBF serving as a strong competitor, as in Section 4.2. Overall, our findings indicate that accurately predicting the performance of FC variants on the complete training dataset based solely on their performance with smaller subsets is challenging. This variability may depend on the specific datasets and the portions of data used for training.

### v4

@@ Abstract
Please refer to the journal version at: https://www.sciencedirect.com/science/article/abs/pii/S0020025526000344
@@ 2.1 KART and KAN
In 1957, Kolmogorov provided a proof to Hilbert’s 13th problem by showing that any multivariate continuous function can be expressed as a combination of single-variable functions and additions, a concept known as KART [26, 6]. This theorem has been utilized in many studies to develop neural networks [63, 42, 27, 34, 31, 30]. However, there is an ongoing debate about the applicability of KART in neural network design. Girosi and Poggio [19] argued that KART’s relevance to neural networks is questionable because the inner function $\phi_{q,p}$ in Equation˜1 may be highly non-smooth [49], which could hinder $f$ from being smooth—a key attribute for generalization and noise resistance in neural networks. Conversely, Kůrková [29] contended that KART is applicable to neural networks, showing that linear combinations of affine functions can effectively approximate all single-variable functions using certain sigmoidal functions.
@@ 3.1 Kolmogorov-Arnold Representation Theorem
A KAN is based on KART, which states that any continuous multivariate function $f$ defined on a bounded domain can be represented as a finite combination of continuous single-variable functions and their additions [8, 40]. For a set of variables $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})\in[0,1]^{n}$, let $f:[0,1]^{n}\to\mathbb{R}$ be a continuous multivariate function. Then $f$ can be expressed as:
@@
which has two types of summations: the outer sum and the inner sum. The outer sum, $\sum_{q=1}^{2n+1}$, aggregates $2n+1$ terms of $\Phi_{q}$ ($\mathbb{R}\to\mathbb{R}$). The inner sum, on the other hand, aggregates $n$ terms for each $q$, where each term $\phi_{q,p}$ ($\phi_{q,p}\colon[0,1]\to\mathbb{R}$) denotes a continuous function of a single variable $x_{p}$.
@@ 3.2 Design of KANs
Remind an MLP that consists of affine transformations and non-linear functions. Starting with an input $\mathbf{x}$, the network processes it through a series of weight matrices across layers (from layer 0 to layer L1L-1) and applies the non-linear activation function $\sigma$ to produce the final output.
@@
Inspired by KART, Liu et al. [37] developed KANs but recommended extending the approach to incorporate greater network widths and depths. To address this, appropriate functions $\Phi_{q}$ and $\phi_{q,p}$ in Equation˜1 need to be identified. A typical KAN network with $L$ layers processes the input $\mathbf{x}$ to produce the output as follows:
@@ 3.3 Implementation of the Current KANs
where $b(x)$ equals to $silu(x)$ as in Equation˜7 and $spline(x)$ is expressed as a linear combination of B-splines $B_{i}$s and control points (coefficients) $c_{i}$s as in Equation˜8. Each activation function is activated with $w_{s}=1$ and $spline(x)\approx 0$, while $w_{b}$ is initialized by using Xavier initialization.
Figure˜1 shows the architecture of KAN(2,3,1), which includes 2 input nodes, 3 hidden nodes, and 1 output node. The output of each node is derived from the sum of individual functions $\phi$, called "edges." The diagram also explains the computation of the inner function $\phi$ using control points (coefficients) and B-splines. The number of B-splines is determined by adding the grid size $G$ and the spline order $k$, resulting in $G+k=3+3=6$, corresponding to the range of $i$ from 0 to 5.
@@
where $w_{i}$ represents adjustable weights or coefficients, and $\phi_{\mathit{RBF}}$ denotes the radial basis function as in Equation˜9.
$\phi_{\mathit{RSWAF}}(r)=1-\left(\tanh\left(\frac{r}{h}\right)\right)^{2}$ (12)
$\mathit{RSWAF}(x)$ $=\sum_{i=1}^{N}w_{i}\phi_{\mathit{RSWAF}}(r_{i})$ (13)
@@
For the simulation of function plots, we present them in Figure˜2, where they are fitted to pass through four selected points. These plots provide a visual representation of the different function types and how they interact with the data points, helping to illustrate their shapes, which are not intended for pre-validating model performance trained on these functions. The functions analyzed include B-spline (3rd-degree), DoG, GRBF, RSWAF, and BSRBF (a combination of B-spline and GRBF). Among these, only the B-spline perfectly passes through all points, while DoG, BSRBF, and GRBF intersect with one point each, and RSWAF passes through none. It is important to note that a function passing through all specified points does not guarantee strong model performance in neural networks, as issues such as overfitting, poor generalization, and sensitivity to outliers may arise.
@@ 3.4 FC-KAN
Figure 4: Various data combinations are performed using element-wise operations (additions ++ and multiplications $\odot$) over two given outputs. The outputs always have the same data dimensions as the inputs.
@@
After that, we combine function outputs to obtain the element-wise output, the concatenated output, and the linearized output, as shown in Figure˜3. The element-wise output is formed by using element-wise operations, including the sum output (Equation˜17a), the element-wise product output (Equation˜17b), the sum and element-wise product output (Equation˜17c), the quadratic and cubic function outputs (Equation˜17d and Equation˜17e). The concatenated output is formed by simply concatenating the function outputs together as in Equation˜17f. The linearized output is formed by passing the concatenated output through a linear transformation (with matrix multiplication by $W$ and addition of bias $b$) as in Equation˜17g. The element-wise output and the linearized output will maintain the same data dimension as each element $\mathbf{o_{i}}$ in $O$, except in the concatenated output, where the data size will be $\mathbf{o_{i}}$ multiplied by the number of function outputs being combined.
@@
Figure˜4 presents data combinations over two given function outputs $\mathbf{o_{1}}$ and $\mathbf{o_{2}}$ using element-wise operations. The results are 5 combined outputs: sum, sum + prod, prod, quad, and cubic. Output combinations can utilize higher-degree functions, but these may significantly increase computational complexity, especially in matrix multiplication. Additionally, using more functions results in a larger number of outputs, which can further complicate data combination calculations. To manage this complexity, we prefer to restrict output combinations to quadratic functions involving up to two outputs. For instance, to combine DoG and B-splines at the output, we can use the following quadratic function formula:
@@ 4.1 Training Configuration
As shown in Table˜1, all models contain a network structure of (784, 64, 10), comprising 784 input neurons, 64 hidden neurons, and 10 output neurons corresponding to the 10 output classes (0-9). Due to the function combinations, FC-KAN has the highest number of parameters, while the MLP has the fewest because it only contains linear transformations over data between layers. The models were trained with 25 epochs on MNIST and 35 epochs on Fashion-MNIST. For KAN models, we use grid_size=5, spline_order=3, and num_grids=8. Other hyperparameters are the same in all models, including batch_size=64, learning_rate=1e-3, weight_decay=1e-4, gamma=0.8, optimize=AdamW, and loss=CrossEntropy.
@@ 4.2 Model Performance
Figure˜5 shows the training losses, represented on a logarithmic scale, for MLP and KAN models on the MNIST and Fashion-MNIST datasets. The loss performance of each model was evaluated based on an independent training run. FC-KAN consistently achieves the lowest training losses across both datasets, followed by BSRBF-KAN due to its fast convergence feature. In contrast, FasterKAN records the highest training loss on MNIST, while MLP performs similarly on Fashion-MNIST.
In general, FC-KAN models outperformed others on MNIST and Fashion-MNIST but require more training time due to the quadratic function representation for output combination, as shown in Section˜4.2. This trade-off between training time and model performance is considered reasonable. The best-performing models are FC-KAN (BS+BASE) and FC-KAN (DoG+BS), which achieved validation accuracies of 97.93% on MNIST and 89.99% on Fashion-MNIST, respectively. When calculating the metric values for both datasets, FC-KAN (DoG+BS) slightly surpassed FC-KAN (BS+BASE) and outperformed the other models. However, FC-KAN (BS+BASE) models take between 9.25% and 11.55% less training time.
@@ 4.3 Comparison of Combination Methods
From the results in Section˜4.3, the quadratic function best represents the combination output and outperforms other combinations, although its models require more training time. It is clear that the outputs combined by element-wise operations consistently outperform other methods, demonstrating superior accuracy. This indicates that element-wise combinations are more effective in capturing and integrating relevant features from the data, leading to better performance. Meanwhile, the output combination by concatenation shows the worst results. The linear transformation applied to the concatenated output outperforms concatenation alone, but still achieves only average performance. Similarly, the maximum, minimum, and average outputs do not deliver superior results.
@@ 4.4 Misclassification Analysis
Figure 6: Heatmaps of the misclassified images in the validation set by models over MNIST and Fashion-MNIST. FC-KAN (DoG+BS) is used for comparison and has the fewest total errors compared to other models.
@@ 4.5 Model Performance with Limited Data
Figure 7: The validation accuracy values of the models across various data subsets. While achieving the highest accuracy values with full data training on MNIST, FC-KAN (BS+BASE) shows no significant advantage with lesser training data. Similarly, FC-KAN (DoG+BS) performs optimally only with 1% of the data.
@@
Figure˜7 illustrates the performance of the FC-KAN models across different training subsets and function combinations. In the MNIST dataset, the RBF+BASE combination demonstrates superior performance with 1%, 5%, and 10% of the data; however, the highest performance on the full training data is recorded for the BS+BASE combination. This observation underscores the challenges of predicting the optimal function combination when working with limited training data. In the Fashion-MNIST dataset, DoG+BS achieves the best performance with 1% of the data, while DoG+RBF excels with both 5% and 10%. These results suggest that DoG+BS and DoG+RBF may perform well with the full training dataset. Indeed, DoG+BS consistently outperforms other combinations in full training, with DoG+RBF serving as a strong competitor, as in Section˜4.2. Overall, our findings indicate that accurately predicting the performance of FC variants on the complete training dataset based solely on their performance with smaller subsets is challenging. This variability may depend on the specific datasets and the portions of data used for training.